# Heterogeneity-Aware Cooperative Federated Edge Learning with Adaptive Computation and Communication Compression

### v1

@@ 5.2 Algorithm Description
To tackle the aforementioned challenges, we develop a solution strategy where the devices (Algorithm 2) and the coordinator (Algorithm 3) collaboratively solve Problem P2. Specifically, at the beginning of each $(l,r)$-th round, each edge device first estimates the local unknown learning variables such as $\sigma_{n}^{l,r}$ and $G_{n}^{l,r}$ and collects the computing parameters $\mu_{n}^{l,r}$, $\alpha_{n}^{l,r}$ and communication parameter $\nu_{n}^{l,r}$, and then sends them to the coordinator as shown in Algorithm 2. After receiving the parameters, the optimal solutions $\bm{\rho}^{*}$ and $\bm{\theta}^{*}$ of Problem P2 can be efficiently computed using an iterative algorithm. The overall algorithm is given in Algorithm 3. Specifically, the coordinator first takes averages of parameters $(\sigma_{n}^{l,r})^{2}$, $(G_{n}^{l,r})^{2}$ to estimate $\sigma^{2}$ and $G^{2}$ (line 1-3). Next, the coordinator solves Problem P2 iteratively by dividing it into two sub-problems.
@@
Problems P2.2 is a quadratic programming problem and can be efficiently solved by a quadratic programming solver, e.g., CVX [38]. In each iteration $e$, the coordinator finds the optimal value of $\bm{\theta}_{e+1}$ while keeping $\bm{\rho}_{e}$ fixed (line 6). Then $\bm{\rho}_{e+1}$ is updated with the obtained $\bm{\theta}_{e+1}$ in the preceding step (line 7). This iterative procedure continues until the convergence criteria $\epsilon$ are met or the maximum number of iterations $I_{\max}$ is reached. Since Problems P2.1 and P2.1 are both convex, both of them can be solved by using some known optimization methods, e.g., interior point method [39].
@@
Algorithm 2 Procedure at device $k$
1:  Estimates $(\sigma_{n}^{l,r})^{2}\leftarrow\mathbb{E}[\|\mathbf{g}_{n}(\mathbf{x}_{n}^{l,r})-\nabla F_{n}(\mathbf{x}_{n}^{l,r})\|^{2}]$
2:  Estimates $(G_{n}^{l,r})^{2}\leftarrow\mathbb{E}[\|\nabla F_{n}(\mathbf{x}_{n}^{l,r})\|^{2}]$
3:  Uploads $(\sigma_{n}^{l,r})^{2}$, $(G_{n}^{l,r})^{2}$, computing parameters $\mu_{n}^{l,r}$, $\alpha_{n}^{l,r}$, communication parameter $\nu_{n}^{l,r}$ to the coordinator.
@@ Appendix A Update Rule for HCEF Scheme
$\mathbf{u}^{t+1}=\mathbf{u}^{t}-\frac{\eta}{N}\sum_{n=1}^{N}Q(\mathbf{g}_{n}^{t}).$ (19)
$\mathbf{g}_{n}^{t}=\mathbbm{1}_{n}^{t}\mathbf{g}_{n}(\mathbf{x}_{n}^{t}),$ (20)
@@
$\mathbb{E}[Q(\mathbf{g}_{n}^{t})]=\mathbb{E}[Q(\mathbb{E}_{\mathbbm{1}_{n}^{t}}[\mathbf{g}_{n}^{t}])]=\rho_{n}^{t}\mathbb{E}[Q([\mathbf{g}_{n}(\mathbf{x}_{n}^{t})])]$ (21)
@@ Appendix D Proof of Lemma 1
Taking the total expectation and averaging over all iterations, then we arrive at the final result.
@@ Appendix E Proof of Lemma 2
$=$ $\eta^{2}\left(\sum_{\alpha=0}^{l-1}\mathbb{E}\|\mathbf{Y}_{\alpha}(\mathbf{Z}^{l-\alpha}-\mathbf{A})\|^{2}+\mathbb{E}\|\mathbf{Y}_{l,r}(\mathbf{V}-\mathbf{A})\|^{2}+\mathbb{E}\|\mathbf{Y}_{l,r,s}(\mathbf{I}-\mathbf{A})\|_{\textup{F}}^{2}\right)$ (27)
$+\underbrace{\eta^{2}\sum_{\alpha=0}^{l-1}\sum_{\alpha^{\prime}=0,\alpha^{\prime}\neq\alpha}^{l-1}\mathbb{E}\underbrace{\left<\mathbf{Y}_{\alpha}(\mathbf{Z}^{l-\alpha}-\mathbf{A}),\mathbf{Y}_{\alpha^{\prime}}(\mathbf{Z}^{l-{\alpha^{\prime}}}-\mathbf{A})\right>}_{TR}}_{TR_{0}}+\underbrace{2\eta^{2}\sum_{\alpha=0}^{l-1}\mathbb{E}\left<\mathbf{Y}_{l,r}(\mathbf{V}-\mathbf{A}),\mathbf{Y}_{\alpha}(\mathbf{Z}^{l-\alpha}-\mathbf{A})\right>}_{TR_{1}}$ (28)
@@
$=\eta^{2}\sum_{\alpha=0}^{l-1}\zeta^{l-\alpha}\mathbb{E}\|\mathbf{Y}_{\alpha}\|_{\textup{F}}^{2}\sum_{{\alpha^{\prime}}=0,\alpha^{\prime}\neq\alpha}^{l-1}\zeta^{(l-{\alpha^{\prime}})}+2\eta^{2}\sum_{\alpha=0}^{l-1}\zeta^{l-\alpha}\mathbb{E}\|\mathbf{Y}_{\alpha}\|_{\textup{F}}^{2}$ (30)
$\quad+\eta^{2}\sum_{\alpha=0}^{l}\zeta^{(l-\alpha)}\mathbb{E}\|\mathbf{Y}_{l,r}\|_{\textup{F}}^{2}+\eta^{2}\sum_{\alpha=0}^{l}\zeta^{(l-\alpha)}\mathbb{E}\|\mathbf{Y}_{l,r,s}\|_{\textup{F}}^{2}$ (31)
$\stackrel{{\scriptstyle\textnormal{(g)}}}{{\mathstrut{\leq}}}\eta^{2}\sum_{\alpha=0}^{l-1}\zeta^{l-\alpha}\mathbb{E}\|\mathbf{Y}_{\alpha}\|_{\textup{F}}^{2}\frac{\zeta}{1-\zeta}+2\eta^{2}\sum_{\alpha=0}^{l-1}\zeta^{l-\alpha}\mathbb{E}\|\mathbf{Y}_{\alpha}\|_{\textup{F}}^{2}$ (32)
@@
where (32) uses the following fact:
@@
$\mathbb{E}\|\mathbf{X}_{t}(\mathbf{I}-\mathbf{A})\|_{\textup{F}}^{2}$ $\leq\eta^{2}\sum_{\alpha=0}^{l-1}\mathbb{E}\|\mathbf{Y}_{\alpha}\|_{\textup{F}}^{2}\zeta^{2(l-\alpha)}+\eta^{2}\mathbb{E}\|\mathbf{Y}_{l,r}\|_{\textup{F}}^{2}+\frac{\eta^{2}}{n}\mathbb{E}\|\mathbf{Y}_{l,r,s}\|_{\textup{F}}^{2}$ (34)
$\quad+\eta^{2}\sum_{\alpha=0}^{l-1}\zeta^{l-\alpha}\mathbb{E}\|\mathbf{Y}_{\alpha}\|_{\textup{F}}^{2}\frac{\zeta}{1-\zeta}+2\eta^{2}\sum_{\alpha=0}^{l-1}\zeta^{l-\alpha}\mathbb{E}\|\mathbf{Y}_{\alpha}\|_{\textup{F}}^{2}$ (35)
@@
$\mathbb{E}\|\mathbf{Y}_{\alpha}\|_{\textup{F}}^{2}$ $=\mathbb{E}\|\sum_{c=\alpha q\tau+1}^{(\alpha+1)q\tau}Q(\mathbf{G}_{c})\|_{\textup{F}}^{2}$ (38)
$=\sum_{n=1}^{N}\mathbb{E}\|\sum_{c=\alpha q\tau+1}^{(\alpha+1)q\tau}Q(\mathbf{g}_{k}^{c})\|^{2}$ (39)
@@
$\mathbb{E}\|\mathbf{X}_{t}(\mathbf{I}-\mathbf{A})\|_{\textup{F}}^{2}\leq$ $\eta^{2}q\tau\sum_{\alpha=0}^{l-1}\left(\zeta^{2(l-\alpha)}+2\zeta^{l-\alpha}+\frac{\zeta^{l-\alpha+1}}{1-\zeta}\right)\sum_{n=1}^{N}\sum_{c=\alpha q\tau+1}^{(\alpha+1)q\tau}\mathbb{E}\|Q(\mathbf{g}_{n}^{c})\|^{2}+\eta^{2}r\tau\left(\frac{2-\zeta}{1-\zeta}\right)\sum_{n=1}^{N}\sum_{c=lq\tau+1}^{lq\tau+r\tau}\mathbb{E}\|Q(\mathbf{g}_{n}^{c})\|^{2}$ (43)
@@
$\sum_{r=0}^{q-1}\sum_{s=0}^{\tau-1}\mathbb{E}\|\mathbf{X}_{t}(\mathbf{I}-\mathbf{A})\|_{\textup{F}}^{2}$ $\leq\eta^{2}q^{2}\tau^{2}\sum_{\alpha=0}^{l-1}\left(\zeta^{2(l-\alpha)}+2\zeta^{l-\alpha}+\frac{\zeta^{l-\alpha+1}}{1-\zeta}\right)\sum_{n=1}^{N}\sum_{c=\alpha q\tau+1}^{(\alpha+1)q\tau}\mathbb{E}\|Q(\mathbf{g}_{n}^{c})\|^{2}$ (44)
$\quad+\eta^{2}\frac{q(q-1)}{2}\tau^{2}\left(\frac{2-\zeta}{1-\zeta}\right)\sum_{n=1}^{N}\sum_{c=lq\tau+1}^{(l+1)q\tau}\mathbb{E}\|Q(\mathbf{g}_{n}^{c})\|^{2}$ (45)
@@
$\sum_{t=0}^{T-1}\mathbb{E}\|\mathbf{X}_{t}(\mathbf{I}-\mathbf{A})\|_{\textup{F}}^{2}\leq\eta^{2}q^{2}\tau^{2}\Omega_{1}\sum_{t=0}^{T-1}\sum_{n=1}^{N}\mathbb{E}\|Q(\mathbf{g}_{n}^{t})\|^{2}.$ (49)
@@ Appendix F Proof of Theorem 1
$+\frac{4G^{2}}{NT}\sum_{t=0}^{T-1}\sum_{n=1}^{N}\left[3(1-\rho_{n}^{t})^{2}+4L^{2}\eta^{2}q^{2}\tau^{2}\Omega_{1}(2-\theta_{n}^{t})\rho_{n}^{t}\right].$ (51)

### v3

@@ 5.2 Algorithm Description
To tackle the aforementioned challenges, we develop a solution strategy where the devices (Algorithm 2) and the coordinator (Algorithm 3) collaboratively solve Problem P2. Specifically, at the beginning of each $(l,r)$-th round, each edge device $n\in\mathcal{S}_{i}$ first estimates the local unknown learning variables such as $\sigma_{n}^{l,r}$ and $G_{n}^{l,r}$ by the received edge model $\mathbf{y}_{i}^{l,r}$. Then, it estimates the computing parameters $\mu_{n}^{l,r}$, $\alpha_{n}^{l,r}$ and communication parameter $\nu_{n}^{l,r}$ according to the batch size, available CPU frequency and channel conditions between the device and server using models in [30, 37, 38] without performing actual training. These parameters are then sent to the coordinator as shown in Algorithm 2. After receiving the parameters, the optimal solutions $\bm{\rho}^{*}$ and $\bm{\theta}^{*}$ of Problem P2 can be efficiently computed using an iterative algorithm. The overall algorithm is given in Algorithm 3. Specifically, the coordinator first takes averages of parameters $(\sigma_{n}^{l,r})^{2}$, $(G_{n}^{l,r})^{2}$ to estimate $\sigma^{2}$ and $G^{2}$ (line 1-3). Next, the coordinator solves Problem P2 iteratively by dividing it into two sub-problems.
@@
Problems P2.2 is a quadratic programming problem and can be efficiently solved by a quadratic programming solver, e.g., CVX [39]. In each iteration $e$, the coordinator finds the optimal value of $\bm{\theta}_{e+1}$ while keeping $\bm{\rho}_{e}$ fixed (line 6). Then $\bm{\rho}_{e+1}$ is updated with the obtained $\bm{\theta}_{e+1}$ in the preceding step (line 7). This iterative procedure continues until the convergence criteria $\epsilon$ are met or the maximum number of iterations $I_{\max}$ is reached. Since Problems P2.1 and P2.1 are both convex, both of them can be solved by using some known optimization methods, e.g., interior point method.
@@
Algorithm 2 Procedure at device $n$
1:  Estimate $(\sigma_{n}^{l,r})^{2}\leftarrow\mathbb{E}[\|\mathbf{g}_{n}(\mathbf{y}_{i}^{l,r})-\nabla F_{n}(\mathbf{y}_{i}^{l,r})\|^{2}]$
2:  Estimate $(G_{n}^{l,r})^{2}\leftarrow\mathbb{E}[\|\nabla F_{n}(\mathbf{y}_{i}^{l,r})\|^{2}]$
3:  Estimates $\mu_{n}^{l,r}$, $\alpha_{n}^{l,r}$, $\nu_{n}^{l,r}$
4:  Uploads $(\sigma_{n}^{l,r})^{2}$, $(G_{n}^{l,r})^{2}$, computing parameters $\mu_{n}^{l,r}$, $\alpha_{n}^{l,r}$, communication parameter $\nu_{n}^{l,r}$ to the coordinator.
@@ 6.3 Ablation Studies
Training parameters $\sigma^{2}$ and $G^{2}$. Fig. 8 illustrates the estimation of the training parameters $\sigma^{2}$ and $G^{2}$ during the execution of the proposed HCEF algorithm. $\sigma^{2}$ represents the variance of the local mini-batch stochastic gradient, while $G^{2}$ tracks the gradient magnitude across. The decreasing trend of $G^{2}$ indicates that our algorithm converges efficiently.
Figure 8: Training parameters $\sigma^{2}$ and $G^{2}$ versus training round.
@@ 7 Conclusions
8 ACKNOWLEDGMENTS
The work of Z. Zhang, Z. Gao, and Y. Gong was partially supported by the US National Science Foundation under Grant CNS-2047761 and Grant CNS-2106761. The work of Y. Guo was partially supported by the US National Science Foundation under Grant CNS-CNS-2106761 and CNS-2318683.
@@ Appendix A Update Rule for HCEF Scheme
$\mathbf{u}^{t+1}=\mathbf{u}^{t}-\frac{\eta}{N}\sum_{n=1}^{N}Q(\mathbf{g}_{n}^{t}).$
$\mathbf{g}_{n}^{t}=\mathbbm{1}_{n}^{t}\mathbf{g}_{n}(\mathbf{x}_{n}^{t}),$
@@
$\mathbb{E}[Q(\mathbf{g}_{n}^{t})]=\mathbb{E}[Q(\mathbb{E}_{\mathbbm{1}_{n}^{t}}[\mathbf{g}_{n}^{t}])]=\rho_{n}^{t}\mathbb{E}[Q([\mathbf{g}_{n}(\mathbf{x}_{n}^{t})])]$
@@ Appendix D Proof of Lemma 1
Taking the total expectation and averaging over all iterations, then we arrive at the final result. ∎
@@ Appendix E Proof of Lemma 2
$=$ $\eta^{2}\left(\sum_{\alpha=0}^{l-1}\mathbb{E}\|\mathbf{Y}_{\alpha}(\mathbf{Z}^{l-\alpha}-\mathbf{A})\|^{2}+\mathbb{E}\|\mathbf{Y}_{l,r}(\mathbf{V}-\mathbf{A})\|^{2}+\mathbb{E}\|\mathbf{Y}_{l,r,s}(\mathbf{I}-\mathbf{A})\|_{\textup{F}}^{2}\right)$
$+\underbrace{\eta^{2}\sum_{\alpha=0}^{l-1}\sum_{\alpha^{\prime}=0,\alpha^{\prime}\neq\alpha}^{l-1}\mathbb{E}\underbrace{\left<\mathbf{Y}_{\alpha}(\mathbf{Z}^{l-\alpha}-\mathbf{A}),\mathbf{Y}_{\alpha^{\prime}}(\mathbf{Z}^{l-{\alpha^{\prime}}}-\mathbf{A})\right>}_{TR}}_{TR_{0}}+\underbrace{2\eta^{2}\sum_{\alpha=0}^{l-1}\mathbb{E}\left<\mathbf{Y}_{l,r}(\mathbf{V}-\mathbf{A}),\mathbf{Y}_{\alpha}(\mathbf{Z}^{l-\alpha}-\mathbf{A})\right>}_{TR_{1}}$
@@
$=\eta^{2}\sum_{\alpha=0}^{l-1}\zeta^{l-\alpha}\mathbb{E}\|\mathbf{Y}_{\alpha}\|_{\textup{F}}^{2}\sum_{{\alpha^{\prime}}=0,\alpha^{\prime}\neq\alpha}^{l-1}\zeta^{(l-{\alpha^{\prime}})}+2\eta^{2}\sum_{\alpha=0}^{l-1}\zeta^{l-\alpha}\mathbb{E}\|\mathbf{Y}_{\alpha}\|_{\textup{F}}^{2}$
$\quad+\eta^{2}\sum_{\alpha=0}^{l}\zeta^{(l-\alpha)}\mathbb{E}\|\mathbf{Y}_{l,r}\|_{\textup{F}}^{2}+\eta^{2}\sum_{\alpha=0}^{l}\zeta^{(l-\alpha)}\mathbb{E}\|\mathbf{Y}_{l,r,s}\|_{\textup{F}}^{2}$
$\stackrel{{\scriptstyle\textnormal{(g)}}}{{\mathstrut{\leq}}}\eta^{2}\sum_{\alpha=0}^{l-1}\zeta^{l-\alpha}\mathbb{E}\|\mathbf{Y}_{\alpha}\|_{\textup{F}}^{2}\frac{\zeta}{1-\zeta}+2\eta^{2}\sum_{\alpha=0}^{l-1}\zeta^{l-\alpha}\mathbb{E}\|\mathbf{Y}_{\alpha}\|_{\textup{F}}^{2}$
@@
where (E) uses the following fact:
@@
$\mathbb{E}\|\mathbf{X}_{t}(\mathbf{I}-\mathbf{A})\|_{\textup{F}}^{2}$ $\leq\eta^{2}\sum_{\alpha=0}^{l-1}\mathbb{E}\|\mathbf{Y}_{\alpha}\|_{\textup{F}}^{2}\zeta^{2(l-\alpha)}+\eta^{2}\mathbb{E}\|\mathbf{Y}_{l,r}\|_{\textup{F}}^{2}+\frac{\eta^{2}}{n}\mathbb{E}\|\mathbf{Y}_{l,r,s}\|_{\textup{F}}^{2}$
$\quad+\eta^{2}\sum_{\alpha=0}^{l-1}\zeta^{l-\alpha}\mathbb{E}\|\mathbf{Y}_{\alpha}\|_{\textup{F}}^{2}\frac{\zeta}{1-\zeta}+2\eta^{2}\sum_{\alpha=0}^{l-1}\zeta^{l-\alpha}\mathbb{E}\|\mathbf{Y}_{\alpha}\|_{\textup{F}}^{2}$
$\quad+\eta^{2}\mathbb{E}\|\mathbf{Y}_{l,r}\|_{\textup{F}}^{2}\frac{1}{1-\zeta}+\eta^{2}\mathbb{E}\|\mathbf{Y}_{l,r,s}\|_{\textup{F}}^{2}\frac{1}{1-\zeta}$
@@
$\mathbb{E}\|\mathbf{Y}_{\alpha}\|_{\textup{F}}^{2}$ $=\mathbb{E}\|\sum_{c=\alpha q\tau+1}^{(\alpha+1)q\tau}Q(\mathbf{G}_{c})\|_{\textup{F}}^{2}$
$=\sum_{n=1}^{N}\mathbb{E}\|\sum_{c=\alpha q\tau+1}^{(\alpha+1)q\tau}Q(\mathbf{g}_{k}^{c})\|^{2}$
@@
$\mathbb{E}\|\mathbf{X}_{t}(\mathbf{I}-\mathbf{A})\|_{\textup{F}}^{2}\leq$ $\eta^{2}q\tau\sum_{\alpha=0}^{l-1}\left(\zeta^{2(l-\alpha)}+2\zeta^{l-\alpha}+\frac{\zeta^{l-\alpha+1}}{1-\zeta}\right)\sum_{n=1}^{N}\sum_{c=\alpha q\tau+1}^{(\alpha+1)q\tau}\mathbb{E}\|Q(\mathbf{g}_{n}^{c})\|^{2}+\eta^{2}r\tau\left(\frac{2-\zeta}{1-\zeta}\right)\sum_{n=1}^{N}\sum_{c=lq\tau+1}^{lq\tau+r\tau}\mathbb{E}\|Q(\mathbf{g}_{n}^{c})\|^{2}$
@@
$\sum_{r=0}^{q-1}\sum_{s=0}^{\tau-1}\mathbb{E}\|\mathbf{X}_{t}(\mathbf{I}-\mathbf{A})\|_{\textup{F}}^{2}$ $\leq\eta^{2}q^{2}\tau^{2}\sum_{\alpha=0}^{l-1}\left(\zeta^{2(l-\alpha)}+2\zeta^{l-\alpha}+\frac{\zeta^{l-\alpha+1}}{1-\zeta}\right)\sum_{n=1}^{N}\sum_{c=\alpha q\tau+1}^{(\alpha+1)q\tau}\mathbb{E}\|Q(\mathbf{g}_{n}^{c})\|^{2}$
$\quad+\eta^{2}\frac{q(q-1)}{2}\tau^{2}\left(\frac{2-\zeta}{1-\zeta}\right)\sum_{n=1}^{N}\sum_{c=lq\tau+1}^{(l+1)q\tau}\mathbb{E}\|Q(\mathbf{g}_{n}^{c})\|^{2}$
@@
$\sum_{t=0}^{T-1}\mathbb{E}\|\mathbf{X}_{t}(\mathbf{I}-\mathbf{A})\|_{\textup{F}}^{2}\leq\eta^{2}q^{2}\tau^{2}\Omega_{1}\sum_{t=0}^{T-1}\sum_{n=1}^{N}\mathbb{E}\|Q(\mathbf{g}_{n}^{t})\|^{2}.$
@@ Appendix F Proof of Theorem 1
$+\frac{4G^{2}}{NT}\sum_{t=0}^{T-1}\sum_{n=1}^{N}\left[3(1-\rho_{n}^{t})^{2}+4L^{2}\eta^{2}q^{2}\tau^{2}\Omega_{1}(2-\theta_{n}^{t})\rho_{n}^{t}\right].$